45 Rationalize the denominator: 5 - 10 The simplified result is
Answer #1

2 45 5-10 =_45(5+v10) (5-110)(5+110) 45 (5+10 (5)110) 45(5 + 110 (25-10) 45 (5+110) 15 = 315+10) = (15+ 3110) ... The simplif